OpenStudy (anonymous):

find the slant height of a cone with surface area 103.67ft^2 and radius 3ft

OpenStudy (anonymous):

is it 8

OpenStudy (kropot72):

The formula for the curved surface of a cone is \[A=\pi \times r \times l\] \[l=\frac{A}{\pi \times r}=\frac{103.67}{\pi \times 3}\]

OpenStudy (kropot72):

@babymambo13 Can you find the slant height now?

OpenStudy (anonymous):

i got 11.005307

OpenStudy (kropot72):

Your answer is correct if you use 3.14 as the value of pi.

OpenStudy (anonymous):

its not one of the answeers

OpenStudy (kropot72):

I expect that 11 is one of the answer choices.

OpenStudy (anonymous):

no the answers they have is 6ft, 9.72ft, 19.44ft, 8ft

OpenStudy (kropot72):

So to calculate the slant height we need to find the lateral area by subtracting the area of the base from the total area. \[lateral\ area=103.67-(\pi \times r ^{2})=75.3957\] \[l=\frac{A}{\pi \times 3}=\frac{75.3957}{\pi \times 3}\]

OpenStudy (anonymous):

its 8 ft

OpenStudy (anonymous):

i knew it

OpenStudy (anonymous):

Thank you

OpenStudy (kropot72):

Yes, your original choice was correct :) You're welcome :)